find the radius of the circle which passes through the origin and the points (a,0) and (0,b)

Grade:11

2 Answers

Arun
6 years ago
Dear student
 
let the equation of circle :
x2 + y2 + 2gx+ 2fy + c = 0
now pass it from (0, 0):
c= 0
 
now pass it from (a, 0)
a2 + 2ag = 0
g = – a/2
 
now pass it from (0, b)
b2 + 2fb = 0
f = – b/2
 
now radius of circle = \sqrt(g2 + f2 –c)
 = \sqrt(a2 + b2)/4
 = ½ \sqrt(a2 + b2)
 
Regards
Arun (askIITians forum expert)
Meet
6 years ago
Let the equation of the circle is x^2+y^2+2gx+2fy+c=0 so if it passes through the origin then c=0, if it passes through (a,0) then a^2+2ga=0 , if passes through (0,b) then b^2+2bf=0 so the equation of circle is x^2+y^2-ax-by=0 so therefore radius of the circle is √(a^2+b^2)/4